OpenStudy (anonymous):

write the equation of a line that is parallel to 4x=3y=7 and also passes through the point (-5,2)

OpenStudy (anonymous):

4x+3y=7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you need to find the slope of that line 3y=-4x+7 y=-(4/3)x+7/3 then you have a point and the slope of the new line, so using point slope form, y-2=-(4/3)(x+5)

OpenStudy (anonymous):

im very lost after u find the slope

OpenStudy (anonymous):

you dont know the slope when the equation of the line is in standard form, so you have to change it to a different form, the easiest form being slope intercept. so to find slope intercept form you solve for y, which is what i did. slope intercept form is y=mx+b where m is the slope. so -(4/3) is the slope. the using the point given, you can find the equation of the parallel line because it will have the same slope.

OpenStudy (anonymous):

then i just took the slope and point and used them to figure out the new equation. point slope form is \[y-y_{1}^{}=m(x-x _{1}^{})\]so y1=2 and x1=-5

OpenStudy (anonymous):

and m = -(4/3)